OpenStudy (anonymous):

Solve the linear inequality: 5 - 5/3x (greater than equal to sign inserted) 9 Sorry I cannot insert the greater than equal to sign into the equation.

OpenStudy (campbell_st):

so you have \[5 - \frac{5}{3x} \ge 9\] start by subtracting 5 from both sides them multiply both sides by 3x this should make it easier for you to solve

OpenStudy (mertsj):

If you multiply both sides by 3x how do you know if you are multiplying by a negative or positive number? Should you switch the inequality symbol or not?

OpenStudy (campbell_st):

well you would have \[-5 \ge 12x\] I hope that make sense...

OpenStudy (anonymous):

Campbell, that is what I got when I multiplied, so would that be the solution set?

OpenStudy (campbell_st):

well there are lots of methods from here... the easiest to understand is to now 1. subtract 12x from both sides 2. add 5 to both sides.. now the inequality will read from left to right as we are normally use to do you know about the inequality and dividing by a negative..?

OpenStudy (anonymous):

No, I do not I have not done inequalities in forever, which is why I am so confused.

OpenStudy (anonymous):

I am confused, instead of adding or subtracting wouldn't you divide -5 in 12x? If I am wrong please let me know, as I said I haven't worked on inequalities in so long

OpenStudy (campbell_st):

well you would have \[-12x \ge 5\] dividing by a negative requires you to reverse the inequality so divide by -12

OpenStudy (mertsj):

\[5-\frac{5}{3x}\ge9\] \[\frac{-5}{3x}-4\ge0\] \[\frac{-5-12x}{3x}\ge0\]

OpenStudy (mertsj):

The numerator is 0 if x = -5/12 The denominator is 0 if x = 0

OpenStudy (anonymous):

Campbell wouldn't that give you just -5/12 ? Oh man now I am really confused.

OpenStudy (campbell_st):

so \[x \le -\frac{5}{12}\]

OpenStudy (anonymous):

Thats what I got, because you need to reserve the sign as well right? I think I am getting this, at least I hope

OpenStudy (mertsj):